OpenStudy (anonymous):

I think its C Which of the following lines is not parallel to the graph of y = 8x + 3? A. 24x – 3y = 4 B. 8x – y = 4 C. y – 8x = 5 D. 8y –x = 4

OpenStudy (anonymous):

nope y – 8x = 5, y = 8x + 5 both r parallel

OpenStudy (anonymous):

Hint: which one has a different slope?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Work it out. :)

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

\[y={-24x+4 \over -3}\]

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

I'll help you even more... y=mx+b where m is the slope. Find the coefficient of x that is different from the rest.

OpenStudy (anonymous):

@monokerous ok so it's B because 8x-y=4 then -y=-8x+4 then y=8x-4

OpenStudy (anonymous):

No... when you bring the negative from y over to -8x+4, the negative becomes positive, right?

OpenStudy (anonymous):

This is B. \[y=-(-8x+4)\]

OpenStudy (anonymous):

From the list of answers, which x doesn't have the coefficient of 8?